The mean number of minutes that android phone users spend on an app is 8.2 minutes. Suppose the standard deviation is 1 minute. Take a sample of 60 What are the mean, standard deviation for the sample mean number of app engagement by a tablet user? and Find the probability that the sample mean is between 8 and 8.5 minutes. 2.) In a recent study, the mean age of online job search website users is 34 years. Suppose the standard deviation is 15 years. Take a sample of size n=100. a.) What are the mean, and standard deviation for the sample mean ages of tablet users? Find the probability that the sample mean age is more than 30 years


1
Expert's answer
2022-05-22T17:52:03-0400

1)

a)


"\mu_{\bar{X}}=\mu=8.2"

"\sigma_{\bar{X}}=\sigma/\sqrt{n}=1/\sqrt{60}\approx0.1291"

b)


"P(8<\bar{X}<8.5)=P(Z<\dfrac{8.5-8.2}{1/\sqrt{60}})"

"-P(Z\le\dfrac{8-8.2}{1/\sqrt{60}})"

"\approx P(Z<2.3238)-P(Z\le-1.5492)"

"\approx0.98993-0.06067\approx0.9293"

2)

a)


"\mu_{\bar{X}}=\mu=34"

"\sigma_{\bar{X}}=\sigma/\sqrt{n}=15/\sqrt{100}=1.5"

b)


"P(\bar{X}>30)=1-P(Z\le\dfrac{30-34}{15/\sqrt{100}})"

"\approx1-P(Z\le-2.66667)"

"\approx0.99617"